# Community-Based Early-Stage Chronic Kidney Disease Screening using Explainable Machine Learning for Low-Resource Settings

### v1

@@ 3.2 Performance of ML Models
Feature subsets derived through feature selection methods (for both S1 and S2) also exhibit strong predictive performance. Across both settings, RFECV-based feature selection consistently outperforms LASSO and the MWU test, highlighting the advantage of model-driven wrapper approaches. Importantly, the selected feature set from S2, which excludes pathology-related variables, achieves performance close to that of the full feature set. Using a CatBoost model on the RFECV+GB-selected S2 feature subset (Table 1), we obtain a balanced accuracy of 89.23%, sensitivity of 83.11%, F1 score of 87.12%, and precision of 90.92%. These values are only 1–2% lower than the best-performing S1 models, demonstrating that CKD can be predicted with competitive accuracy using solely non-pathological, non-laboratory features. Figures 2 and 3 together illustrate the comparative performance and classification behaviour of models trained on those feature sets. This finding underscores the feasibility of deploying CKD screening models in community-based and low-resource settings without the need for laboratory or pathology tests.
@@
BalancedAccuracySensitivity (CKD)AUCROCF1 (CKD)Precision0.60.60.70.70.80.80.90.911Performance scoreAllBest (S2)Best (S1)
@@
Figure 2: Performance comparison of machine learning models trained on three different feature configurations: the full feature set, the best-performing subset excluding pathology features (S2), and the best-performing subset across all feature domains (S1). Error bars represent 95% confidence intervals computed over 10-fold cross validation.
@@
3.3 Comparison with State-Of-The-Art Tools
Table 4 provides a comprehensive comparison between the proposed CKD prediction models and several state-of-the-art (SOTA) CKD screening tools, including SCORED [bang2007development], kshirsagar2008simple (simplified version), thakkinstian2011simple, kearns2013predicting, and kwon2019simple. For each tool, the table reports balanced accuracy, sensitivity (CKD), AUC-ROC, F1 score (CKD), and macro-precision. The two proposed models, one using the best-performing S1 feature set with a Decision Tree, and the other using the best S2 subset with CatBoost, demonstrate substantially higher balanced accuracy, sensitivity, F1 score, and precision than all existing tools.
Table 4: Performance comparison with state-of-the-art (SOTA) CKD screening tools. The SPS tool does not directly classify individuals as CKD or non-CKD; instead, it assigns four risk categories: low, intermediate-low, intermediate-high, and high. Accordingly, the low category was treated as non-CKD, while the remaining three categories were treated as CKD for evaluation. Asterisks following each performance value indicate the results of pairwise statistical comparisons between the proposed model and the baseline tools.
Tool Balanced accuracy Sensitivity (CKD) AUC-ROC F1 (CKD) Precision (macro)
SCORED [bang2007development] 76.35(∗∗∗S1,∗∗∗S2) 66.07(∗∗∗S1,∗∗∗S2) 0.8507(∗S1,∗∗∗S2) 70.81(∗∗∗S1,∗∗∗S2) 77.98(∗∗S1,∗∗∗S2)
[…]
kwon2019simple 77.49(∗∗∗S1,∗∗∗S2) 64.29(∗∗∗S1,∗∗∗S2) 0.8537(∗S1,∗∗S2) 72.00(∗∗∗S1,∗∗∗S2) 80.71(∗∗S1,∗∗∗S2)
[…]
This study (best S1 with DT) 90.40 84.85 0.9062 88.77 92.14
This study (best S2 with CB) 89.23 83.11 0.9062 87.12 90.92
*$p<0.05$, **$p<0.01$, ***$p<0.001$
Asterisks following each performance value indicate the statistical significance of improvements relative to the SOTA methods, with comparisons presented in a fixed order (S1, S2). Most comparisons show highly significant (**$p<0.01$) or extremely significant (***$p<0.001$) improvements, underscoring that the proposed models offer markedly superior predictive performance over current screening tools. Overall, the table highlights the strong advantage of our feature selection strategy and machine learning framework in achieving more accurate and clinically meaningful CKD screening.
@@ 3.4 External Validation
To evaluate the applicability and generalizability of the best-performing S1 and S2 feature subsets, we conducted external validation using three independent publicly available datasets: UCI-2023 [chronic_kidney_disease_336], UCI-2015 [risk_factor_prediction_of_chronic_kidney_disease_857], and TH [Al-Shamsi2018]. For each of these datasets, we first identified the set of features shared with our dataset and harmonized the naming conventions to ensure consistency across sources. Based on these harmonized features, we then constructed the S1 and S2 subsets by selecting the variables that overlapped with our optimized S1 and S2 feature sets, respectively. The resulting common, S1, and S2 feature subsets for each external dataset are summarized in Table 5. Notably, the TH dataset did not contain any pathology-related features that matched those in our dataset, and therefore, its S1 subset could not be constructed. This process enables a direct and fair assessment of how well the models trained on our dataset generalize when applied to diverse populations and measurement contexts.
@@
UCI-2023 [chronic_kidney_disease_336] Common Hypertension, Age_60+y, Diabetes, Anemia, RBC, Age_31-39y
@@
S2 subset Hypertension, Age_60+y, Diabetes, Anemia, Age_31-39y
UCI-2015 [risk_factor_prediction_of_chronic_kidney_disease_857] Common Hypertension, Age_60+y, Diabetes, Anemia, RBC, Age_18-30y,
@@
S2 subset Hypertension, Age_60+y, Diabetes, Anemia, Age_31-39y
TH [Al-Shamsi2018] Common Hypertension, Age_60+y, Diabetes, BMI_Obese, Gender, Age_31-39y,
@@
S2 subset Hypertension, Age_60+y, Diabetes, Age_31-39y
Table 6 presents the external validation performance of the machine learning models trained on our dataset and evaluated on the three independent datasets. Across all datasets, both the common and subset-based models demonstrate strong sensitivity (e.g., 80.47% in UCI-2023) and competitive balanced accuracy (e.g., 79.33% in UCI-2015), indicating good generalizability of the trained models to external populations. Notably, the TH dataset contains only CKD cases; therefore, only sensitivity is reported for this cohort, with the S2 subset demonstrating exceptional performance at 98.21%.
Table 6: External validation performance of machine learning models trained on our dataset and tested on three external datasets.
Test dataset Feature set Balanced accuracy Sensitivity (CKD) AUC-ROC F1 (CKD) Precision (macro)
UCI-2023 [chronic_kidney_disease_336] Common 76.35 80.47 0.819 82.07 75.64
S1 subset 76.35 80.47 0.844 82.07 75.64
S2 subset 74 75.78 0.804 79.18 72.78
UCI-2015 [risk_factor_prediction_of_chronic_kidney_disease_857] Common 78.93 77.2 0.826 81.78 77.46
S1 subset 79.33 78 0.881 82.28 77.9
S2 subset 77.13 73.6 0.842 79.48 75.55
TH [Al-Shamsi2018] Common 96.43
S2 subset 98.21
Across all datasets, models using the S1 subset (for UCI-2023 and UCI-2015) and the S2 subset (for TH) consistently outperform those using the common feature set, demonstrating improved discriminatory power. This consistent improvement underscores the effectiveness of the feature selection strategy developed in this study and supports its generalizability and applicability across diverse datasets.
@@ 4.1 Insights and Observations
Second, the results reaffirm that adding more features does not necessarily improve classification performance. For S1, the common feature subset (Hypertension, Age_60+y, and RBC) outperforms both the union and all-feature sets, despite the latter containing considerably more variables. Similarly, for S2, the common subset (Hypertension, Age_60+y) surpasses the performance of the largest S2 feature set (SD–LH–MH–CE). These findings emphasize that a compact set of well-chosen, clinically relevant features can outperform larger, noisier combinations. This has direct implications for model interpretability, ease of data collection, and feasibility of large-scale screening in low-resource environments.
Third, the common feature subsets themselves deliver remarkably strong performance. In S1, the three-feature subset (Hypertension, Age_60+y, RBC) achieves a balanced accuracy of 86.19%, which is highly competitive and sufficient for clinical triage despite its minimal size. Even more striking is the S2 common subset (Hypertension and Age_60+y), which achieves 83.50% balanced accuracy while requiring only non-pathological, easily obtainable features. This highlights the significant potential for deploying CKD risk assessment tools in settings lacking laboratory infrastructure, where pathology tests may be inaccessible.
@@
Fourth, while no single classifier consistently dominates across all feature sets, we observe that many models achieve closely comparable performance, reflecting the stable predictive structure in the data. However, AdaBoost emerges as the most frequently top-performing classifier (Figure 1). This suggests that boosting-based ensemble methods may be particularly well-suited for integrating heterogeneous, mixed-type clinical features when predicting CKD.
@@
Fifth, although our focus was on early-stage CKD detection (stages 1–3), the model demonstrates excellent performance in detecting advanced-stage CKD as well. The TH dataset, which comprises only CKD stages 3–5, achieves an exceptional sensitivity of 98.21% using the S2 feature subset. This robustness across disease stages suggests that the selected features capture fundamental CKD risk patterns regardless of severity. Moreover, successful external validation on datasets from India (UCI-2015) and the UAE (TH) indicates that our approach generalizes reasonably well across ethnically diverse populations in South Asia and the Middle East.
@@
Finally, the comparison with state-of-the-art CKD screening tools reveals that our models not only outperform existing methods in accuracy, sensitivity, and AUC-ROC but also rely on fewer and more readily available features. Tools such as SCORED, kshirsagar2008simple, kshirsagar2008simple, and kearns2013predicting often require broader or more specialized input variables (see Table 2), some of which may not be accessible in low-resource settings. In contrast, our optimized S1 and S2 subsets achieve superior performance with simpler, more scalable feature requirements, highlighting the practical advantage of our methodology for population-level CKD screening.
4.2 Limitations and Future Work
Although this study demonstrates strong predictive performance and compelling generalizability for early CKD detection, several limitations should be acknowledged.
First, the primary dataset used for model development was collected from a specific geographic region and demographic profile, which may not fully represent broader global populations. While external validation using datasets from India and the UAE supports cross-population generalizability, further evaluation on datasets from Africa, Europe, and Latin America would strengthen evidence of worldwide applicability.
Second, the dataset size was relatively small; however, this is the first CKD screening study conducted following the standard CKD diagnosis technique in Bangladesh and, more broadly, across South Asia. Although large datasets such as the UK Biobank or the U.S. CRIC cohort exist, they represent substantially different ethnic groups and are therefore unlikely to reflect South Asian patient characteristics. Moreover, such datasets are not readily accessible. The three external datasets we used also lacked complete alignment with our feature space and CKD staging structure. For example, none of them included the full set of best-performing features identified in our study, limiting the ability to perform comprehensive cross-validation. Additionally, the TH dataset contained only CKD stage 3–5 cases and no non-CKD participants, restricting the calculation of performance metrics that require both classes (e.g., balanced accuracy, AUC). These constraints highlight the urgent need for more complete, high-quality CKD datasets, particularly in underrepresented regions, to support broader validation.
Third, the cross-sectional nature of the datasets prevents evaluation of model performance in predicting CKD progression or transitions between disease stages. Longitudinal datasets are essential for developing models that not only detect CKD early but also forecast disease trajectories and inform personalized intervention strategies.
Fourth, although the primary dataset was imbalanced, we elected not to apply oversampling techniques, as these would introduce synthetic data that may not reflect true clinical patterns. Instead, we employed the “balanced” class-weight parameters available in several machine learning algorithms. While this preserves data authenticity, future studies may explore advanced imbalance-handling methods that avoid synthetic data generation, such as focal loss or cost-sensitive learning.
Fifth, despite evaluating a comprehensive set of machine learning classifiers, our analysis focused on classical ML methods. Although deep learning does not always outperform classical approaches for tabular clinical data, future work could examine hybrid neural models, attention-based architectures, or representation-learning strategies capable of capturing more complex feature interactions.
@@ 5 Conclusion
This study presents a comprehensive and data-driven framework for early CKD detection using machine learning, demonstrating that high predictive performance can be achieved with minimal, non-invasive, and readily obtainable features. Through rigorous feature selection, extensive model evaluation, comparison with state-of-the-art tools, and validation across three independent external datasets, our findings highlight the robustness, generalizability, and clinical relevance of the proposed approach. Notably, the S2 feature subset, requiring only two non-laboratory variables, achieves performance comparable to full feature sets, underscoring its potential for use in community-based and resource-limited settings where laboratory testing is not feasible. The explainability analysis further enhances transparency and trust by revealing clinically meaningful feature contributions. Overall, this work provides a strong foundation for developing scalable CKD screening solutions and paves the way for future advancements, including longitudinal modeling, integration with digital health platforms, and prospective evaluation in real-world clinical environments.
@@
Of the four datasets used in this study, three are publicly available and can be accessed through their respective repositories as cited in the manuscript. The ICDDR,B dataset cannot be made publicly available due to ethical and institutional restrictions.
Appendix A Dataset summary
Table 1: Summary of CKD datasets used in this study
Dataset Country Sample size CKD Stage Features
CKD Non-CKD Total SD LH MH CE Path Path* Total
UCI-2015 [chronic_kidney_disease_336] India 250 150 400 1 1 3 5 12 2 24
UCI-2023 [risk_factor_prediction_of_chronic_kidney_disease_857] Bangladesh 128 72 200 1–5 1 1 4 6 12 3 27
TH [Al-Shamsi2018] UAE 56 56 3–5 2 1 10 3 3 2 21
This study Bangladesh 112 172 284 1–3 5 3 7 4 5 5 29
*Gold-standard pathology tests for CKD diagnosis, such as serum creatinine, urinary albumin, urinary creatinine, eGFR, and uACR
Appendix B Performance metrics definition

### v2

@@ 4.2 Performance of ML Models
Feature subsets derived through feature selection methods (for both S1 and S2) also exhibit strong predictive performance. Across both settings, RFECV-based feature selection consistently outperforms LASSO and the MWU test, highlighting the advantage of model-driven wrapper approaches. Importantly, the selected feature set from S2, which excludes pathology-related variables, achieves performance close to that of the full feature set. Using a CatBoost model on the RFECV+GB-selected S2 feature subset (Table 6), we obtain a balanced accuracy of 89.23% (95% CI 84.9%–92.6%), sensitivity of 83.11% (95% CI 74.1%–88.4%), F1 score of 87.12% (95% CI 82.0%–91.4%), and precision of 90.92% (95% CI 86.8%–93.7%). These values are only 1–2% lower than the best-performing S1 models, demonstrating that CKD can be predicted with competitive accuracy using solely non-pathological, non-laboratory features. Figures 2 and 3 together illustrate the comparative performance and classification behaviour of models trained on those feature sets. This finding underscores the feasibility of deploying CKD screening models in community-based and low-resource settings without the need for laboratory or pathology tests.
@@
BalancedAccuracySensitivity (CKD)AUCROCF1 (CKD)Precision0.60.60.70.70.80.80.90.911Performance scoreAll featuresBest feature subset w/o pathology (S2)Best feature subset among all (S1)
@@
Figure 2: Performance comparison of machine learning models trained on three different feature configurations: the full feature set, the best-performing subset excluding pathology features (S2), and the best-performing subset across all feature domains (S1). Performance metrics are reported as the average across 10-fold cross-validation, and error bars represent 95% confidence intervals estimated using stratified nonparametric BCa bootstrap resampling of pooled fold predictions from all folds, which may produce asymmetric intervals.
@@
4.3 Comparison with CKD Screening Tools
Table 7 compares the proposed machine learning models with several established CKD screening tools, including SCORED [6], Kshirsagar et al. [46] (simplified version), Thakkinstian et al. [71], Kearns et al. [42], and Kwon et al. [47]. For consistency with the binary classification framework used in this study, the Thakkinstian et al. [71] tool, originally designed to assign individuals to four risk categories (low, intermediate-low, intermediate-high, and high), was converted into a binary prediction task by treating the low category as non-CKD and the remaining categories as CKD. In addition, the probability output of Kearns et al. [42] does not naturally scale between 0 and 1; therefore, an optimal decision threshold was determined by maximizing the F1 score on the evaluation data, resulting in a threshold value of 0.0363. A brief description of these tools is provided in B and summarized in 1.
As shown in Table 7, the proposed models based on the optimized S1 and S2 feature subsets consistently outperform all existing screening tools across most evaluation metrics. To assess whether the observed improvements are statistically significant, we performed a paired permutation test based on per-sample Brier loss using the shared evaluation samples. The resulting $p$-values are reported in Table 7 for both the S1 and S2 models. The results show that the proposed models significantly outperform all existing screening tools ($p<0.001$). These results indicate that the machine learning models derived from the proposed feature selection framework provide substantially improved predictive performance compared with traditional risk-score–based screening tools.
Table 7: Performance comparison between the proposed machine learning models and established CKD screening tools on the TangailBD dataset. S1 denotes the optimized feature subset derived from all variables, and S2 denotes the optimized subset excluding pathology-test variables. The proposed models (best S1 with Decision Tree and best S2 with CatBoost) are evaluated against traditional screening tools. Statistical significance is assessed using a paired permutation test based on per-sample Brier loss, with $p$-values indicating whether the performance differences between the proposed models and each screening tool are statistically significant.
Tool Balanced accuracy Sensitivity (CKD) AUC-ROC F1 (CKD) Precision (macro) Significance
S1 S2
SCORED [6] $76.35_{\text{70.9--81.2}}$ $66.07_{\text{55.3--73.2}}$ $0.851_{\text{0.79--0.89}}$ $0.81_{\text{63.6--77.2}} 77$$7.98_{\text{72.4--82.6}}$ $p<0.001$ $p<0.001$
Kshirsagar et al. [46] $77.09_{\text{71.7--81.8}}$ $66.96_{\text{57.1--74.1}}$ $0.869_{\text{0.82--0.91}}$ $71.77_{\text{64.6--77.9}}$ $78.77_{\text{73.4--83.5}}$ $p<0.001$ $p<0.001$
Thakkinstian et al. [71] $81.03_{\text{75.9--85.4}}$ $81.25_{\text{72.3--86.6}}$ $0.804_{\text{0.74--0.86}}$ $77.12_{\text{71.4--82.2}}$ $80.13_{\text{75.2--84.5}}$ $p<0.001$ $p<0.001$
Kwon et al. [47] $77.49_{\text{72.2--82.2}}$ $64.29_{\text{53.6--71.4}}$ $0.854_{\text{0.80--0.89}}$ $72.00_{\text{64.3--78.5}}$ $80.71_{\text{75.1--85.2}}$ $p<0.001$ $p<0.001$
Kearns et al. [42] $78.73_{\text{73.5--83.1}}$ $83.13_{\text{74.1--88.4}}$ $0.867_{\text{0.81--0.90}}$ $4.70_{\text{69.1--79.5}} 77$$7.48_{\text{72.4--81.7}}$ $p<0.001$ $p<0.001$
This study (best S1 with DT) $90.40_{\text{86.1--93.5}}$ $84.85_{\text{75.9--90.2}}$ $0.907_{\text{0.86--0.94}}$ $88.77_{\text{83.5--92.6}}$ $92.14_{\text{87.9--94.7}}$ $p=0.74$
This study (best S2 with CB) $89.23_{\text{84.9--92.6}}$ $83.11_{\text{74.1--88.4}}$ $0.906_{\text{0.84--0.94}}$ $87.12_{\text{82.0--91.4}}$ $90.92_{\text{86.8--93.7}}$ $p=0.74$
Importantly, the comparison between the S1 and S2 models shows no statistically significant difference ($p=0.74$), indicating that the reduced feature subset excluding pathology-test variables achieves performance comparable to the full feature set. This finding suggests that effective CKD screening can be achieved using minimal non-laboratory features, which are more suitable for community-level screening in resource-constrained settings.
@@ 4.4 External Validation
To evaluate the applicability and generalizability of the best-performing S1 and S2 feature subsets, we conducted external validation using three independent publicly available datasets: UCI-2023 [64], UCI-2015 [35], and TH [4]. For each of these datasets, we first identified the set of features shared with our dataset and harmonized the naming conventions to ensure consistency across sources. Based on these harmonized features, we then constructed the S1 and S2 subsets by selecting the variables that overlapped with our optimized S1 and S2 feature sets, respectively. The resulting common, S1, and S2 feature subsets for each external dataset are summarized in Table 8. Notably, the TH dataset did not contain any pathology-related features that matched those in our dataset, and therefore, its S1 subset could not be constructed. This process enables a direct and fair assessment of how well the models trained on our dataset generalize when applied to diverse populations and measurement contexts.
@@
UCI-2023 [64] Common Hypertension, Age_60+y, Diabetes, Anemia, RBC
@@
S2 subset Hypertension, Age_60+y, Diabetes, Anemia
UCI-2015 [35] Common Hypertension, Age_60+y, Diabetes, Anemia, RBC, Age_18-30y,
@@
S2 subset Hypertension, Age_60+y, Diabetes, Anemia, Age_18-30y
TH [4] Common Hypertension, Age_60+y, Diabetes, BMI_Obese, Gender, Age_31-39y,
@@
S2 subset Hypertension, Age_60+y, Diabetes, Age_18-30y
Table 9 presents the external validation performance of the machine learning models trained on our dataset and evaluated on the three independent datasets. Across all datasets, both the common and subset-based models demonstrate strong sensitivity (e.g., 80.47%, 95% CI 71.9%–85.9%, in UCI-2023) and competitive balanced accuracy (e.g., 79.33%, 95% CI 74.9%–83.1%, in UCI-2015), indicating good generalizability of the trained models to external populations. Notably, the TH dataset contains only CKD cases; therefore, only sensitivity is reported for this cohort, with the S2 subset demonstrating exceptional performance at 98.21% (95% CI 89.3%–100.0%).
Table 9: External validation performance of machine learning models trained on TangailBD dataset and tested on three external datasets.
Train Test Feature set B. Accuracy Sensitivity AUC-ROC F1 Precision (macro)
TangailBD UCI-2023 [64] Common $76.35_{\text{69.5--82.0}}$ $80.47_{\text{71.9--85.9}}$ $0.819_{\text{0.76--0.87}}$ $82.07_{\text{76.8--86.5}}$ $75.64_{\text{68.9--81.4}}$
S1 subset $76.35_{\text{69.5--82.0}}$ $80.47_{\text{71.9--85.9}}$ $0.844_{\text{0.79--0.89}}$ $82.07_{\text{76.8--86.5}}$ $75.64_{\text{68.9--81.4}}$
S2 subset $74.00_{\text{67.1--79.8}}$ $75.78_{\text{66.4--82.0}}$ $0.804_{\text{0.74--0.86}}$ $79.18_{\text{73.4--84.0}}$ $72.78_{\text{66.2--78.3}}$
UCI-2015 [35] Common $8.93_{\text{74.4--82.8}} 77$$7.20_{\text{71.2--81.6}}$ $0.826_{\text{0.78--0.86}}$ $81.78_{\text{77.8--85.2}}$ $77.46_{\text{73.1--81.2}}$
S1 subset $79.33_{\text{74.9--83.1}}$ $78.00_{\text{72.0--82.4}}$ $0.881_{\text{0.84--0.91}}$ $82.28_{\text{78.3--85.7}}$ $77.90_{\text{73.7--81.6}}$
S2 subset $77.13_{\text{72.6--80.9}}$ $73.60_{\text{67.2--78.2}}$ $0.842_{\text{0.79--0.88}}$ $79.48_{\text{75.2--83.1}}$ $75.55_{\text{71.3--79.2}}$
TH [4] Common $96.43_{\text{85.6--98.2}}$
S2 subset $98.21_{\text{89.3--100}}$
B. Accuracy - Balanced Accuracy
@@ 5 Discussion
Second, while no single classifier consistently dominates across all feature sets, we observe that many models achieve closely comparable performance, reflecting the stable predictive structure in the data. However, AdaBoost emerges as the most frequently top-performing classifier (Figure 6). This suggests that boosting-based ensemble methods may be particularly well-suited for integrating heterogeneous, mixed-type clinical features when predicting CKD.
@@
Third, although our focus was on early-stage CKD detection (stages 1–3), the model demonstrates excellent performance in detecting advanced-stage CKD as well. The TH dataset, which comprises only CKD stages 3–5, achieves an exceptional sensitivity of 98.21% using the S2 feature subset. This robustness across disease stages suggests that the selected features capture fundamental CKD risk patterns regardless of severity. Moreover, successful external validation on datasets from India (UCI-2015) and the UAE (TH) indicates that our approach generalizes reasonably well across ethnically diverse populations in South Asia and the Middle East.
@@
Fourth, the comparison with state-of-the-art CKD screening tools reveals that our models not only outperform existing methods in accuracy, sensitivity, and AUC-ROC but also rely on fewer and more readily available features. Tools such as SCORED, Kshirsagar et al. [46], Kshirsagar et al. [46], and Kearns et al. [42] often require broader or more specialized input variables (see Table 1), some of which may not be accessible in low-resource settings. In contrast, our optimized S1 and S2 subsets achieve superior performance with simpler, more scalable feature requirements, highlighting the practical advantage of our methodology for population-level CKD screening.
@@
In this study, individuals with advanced CKD (stage 4–5) were excluded from the primary analytic dataset because the primary objective was to develop models for community-based early-stage CKD screening (stages 1–3). In the TangailBD dataset, this exclusion involved a small number of participants (stage 4: n = 5; stage 5: n = 1). While this design focuses the model on early disease detection, it does not prevent the identification of advanced CKD cases in practice. Individuals with later-stage CKD typically exhibit stronger risk signals and would still be expected to be classified as high-risk. This is supported by the external validation results, where the model achieved high sensitivity on datasets containing more advanced CKD cases. In real-world deployment, the proposed framework is intended as a screening tool to identify individuals at elevated CKD risk, after which confirmatory clinical evaluation and laboratory testing would be used to determine disease stage and guide treatment.
A key limitation of this study is the relatively modest size and geographic scope of the primary dataset (TangailBD), which was collected from a specific region in Bangladesh. While the dataset is community-based and appropriate for the study objective, its limited sample size and regional focus may affect the robustness and representativeness of the learned models. Although external validation was conducted on three independent datasets to assess generalizability, further evaluation on larger and more diverse community-based populations is necessary to confirm the stability and broader applicability of the proposed framework.
In addition, although the proposed models demonstrate strong predictive performance, their effectiveness may be influenced by variations in data distribution and feature availability across populations, highlighting the importance of further validation under diverse real-world conditions.
@@ 6 Conclusion
This study presents a comprehensive and data-driven framework for early CKD detection using machine learning, demonstrating that strong predictive performance can be achieved with minimal, non-invasive, and readily obtainable features. Through rigorous feature selection, extensive model evaluation, comparison with state-of-the-art tools, and validation across three independent external datasets, our findings highlight the robustness, generalizability, and clinical relevance of the proposed approach. Notably, the S2 feature subset, requiring only two non-laboratory variables, achieves performance comparable to full feature sets, underscoring its potential for use in community-based and resource-limited settings where laboratory testing is not feasible. The explainability analysis further enhances transparency and trust by revealing clinically meaningful feature contributions. Overall, this work provides a strong foundation for future development of scalable CKD screening solutions. Further studies evaluating implementation feasibility, cost-effectiveness, and integration within community health systems will be required before real-world deployment can be considered.
@@
Of the four datasets used in this study, three are publicly available and can be accessed through their respective repositories as cited in the manuscript. The ICDDR,B dataset cannot be made publicly available due to ethical and institutional restrictions. Data requests for ICDDR,B can be directed to: habibur.rahman@icddrb.org.
Appendix A Performance metrics definition